# LICO: Large Language Models for In-Context Molecular Optimization

### v1

@@ 5.1 PMO Benchmark
Benchmark We evaluate LICO on Practical Molecular Optimization (PMO) [14], a standard benchmark for molecular optimization with a focus on sample efficiency. We experiment on $21$ optimization objectives provided by PMO, including QED [6], DRD2 [39], and $19$ objective functions from Guacamol [9]. QED assesses a molecule’s drug-likeness by identifying certain "red flags". DRD2 is a machine learning model trained on experimental data to predict bioactivities for specific target diseases. Guacamol objectives emulate drug discovery goals through a multi-property objective (MPO) approach, considering factors like target molecule similarity, molecular weights, and CLogP. All objective values range from 00 to $1$, with $1$ indicating the best outcome.
@@
Baselines We compare LICO against $3$ leading methods in PMO, namely REINVENT [39], Graph GA [21], and GP BO [46]. REINVENT is a reinforcement learning method that refines a pretrained RNN for generating SMILES strings. Graph GA, inspired by evolutionary processes, utilizes crossover operations derived from graph matching and mutation at atom and fragment levels to explore the molecule space. GP BO is a Bayesian optimization method that constructs a surrogate function using Gaussian Processes and employs an acquisition function that combines the surrogate’s predictions with uncertainty estimates to guide candidate selection. In addition to PMO baselines, we also compare LICO with TNP [36], a state-of-the-art transformers model for few-shot learning. GP BO and TNP are similar to LICO, where the difference is we use an LLM for surrogate modeling instead of a GP or TNP.
@@
LICO training We use ZINC 250K as the unlabeled dataset $\mathcal{D}_{\text{u}}$. ZINC 250K contains around $250000$ molecules sampled from the full ZINC database [42] with moderate size and high pharmaceutical relevance and popularity. We adopt $2$-radius $2048$ bit molecular fingerprints as the input feature of the molecule. To generate training data, we use $47$ intrinsic properties of the molecule as the intrinsic functions, which we present in detail in Appendix A.1. We train LICO for $20000$ iterations with a batch size of $4$, where each data point is a sequence of $(x,y)$ pairs sampled from an intrinsic or synthetic function. The ratio of synthetic data is $0.1$. We use Llama-2-7b [45] as the base LLM, and use LoRA [19] for parameter-efficient finetuning.
@@
Optimization details We limit the optimization budget of all methods to $1000$ function calls. We report the area under the curve (AUC) of the top-$10$ average objective value against the number of function calls as the performance metric. AUC metric favors methods that obtain high values with a smaller number of function calls, thus evaluating both optimization capability and sample efficiency. We min-max scale the AUC values to $[0,1]$. We aggregate the performance for each method across $5$ seeds for better reproducibility as suggested by PMO.
@@
Table 1: The performance of LICO and the baselines on $21$ optimization tasks in PMO. Higher score is better. We report the mean and stddev of scores averaged over $5$ random seeds. We use blue and violet to denote the best and second-best method for each task.

### v2

@@ 5.1 PMO Benchmark
[…]
@@
Baselines We compare LICO against $6$ leading methods in PMO, namely Genetic GFN (Kim et al., 2024), REINVENT (Olivecrona et al., 2017), Augmented Memory (Guo & Schwaller, 2023), Graph GA (Jensen, 2019), GP BO (Tripp et al., 2021), and MOLLEO (Wang et al., 2024). Genetic GFN employs a GFlowNets (Bengio et al., 2023) model trained to sample molecules proportional to their rewards. REINVENT is a reinforcement learning method that finetunes a pretrained RNN for generating SMILES strings, and Augmented Memory combines REINVENT with data augmentation and experience replay. Graph GA, inspired by evolutionary processes, utilizes crossover and mutation operations to explore the molecule space. GP BO is a Bayesian optimization method that augments Graph GA with a Gaussian Processes surrogate model and UCB acquisition function to guide candidate selection. MOLLEO is an LLM-based method that prompts a chemistry-aware LLM such as BioT5 (Pei et al., ) to perform mutation and crossover operations in an evolutionary algorithm. Among the baselines, GP BO is the most similar to LICO, where the only difference is we use an LLM for surrogate modeling instead of a GP.
@@
LICO training We use ZINC 250K as the unlabeled dataset $\mathcal{D}_{\text{u}}$. ZINC 250K contains around $250000$ molecules sampled from the full ZINC database (Sterling & Irwin, 2015) with moderate size and high pharmaceutical relevance and popularity. We adopt $2$-radius $2048$ bit Morgan molecular fingerprints as the input feature of the molecule. To generate training data, we use $47$ intrinsic properties of the molecule as the intrinsic functions, which we present in detail in Appendix A.1. We train LICO for $20000$ iterations with a batch size of $4$, where each data point is a sequence of $(x,y)$ pairs sampled from an intrinsic or synthetic function. The ratio of synthetic data is $0.1$. We use Llama-2-7b (Touvron et al., 2023b) as the base LLM, and use LoRA (Hu et al., 2021) for parameter-efficient finetuning. We use the Llama-2-7b-chat checkpoint for the 1000-budget setting, and the Llama-2-7B-32K-Instruct checkpoint with the Liger Kernel (Hsu et al., 2024) for training and inference with long context for the 10000-budget setting.
@@
Evaluation details We report the area under the curve (AUC) of the top-$10$ average objective value against the number of function calls as the performance metric. AUC metric favors methods that obtain high values with a smaller number of function calls, thus evaluating both optimization capability and sample efficiency. We min-max scale the AUC values to $[0,1]$. We aggregate the performance for each method across $5$ seeds for better reproducibility as suggested by PMO.
@@
Table 1: The performance of LICO and the baselines on $23$ optimization tasks in PMO-1K. A higher score is better. We report the mean and stddev of scores averaged over 55 random seeds. We use blue and violet to denote the best and second-best method for each task.